e and a aded Odre e s rued an a Pun ot 5 s oled and let & be eved fhud a umo
Answer #1

The sample space consist of 6 * 4 = 24 sample points as:

{1, 1}; {1, 2}; {1, 3}; {1, 4} ;

{2, 1}; {2, 2}; {2, 3}; {2, 4} ;

{3, 1}; {3, 2}; {3, 3}; {3, 4} ;

{4, 1}; {4, 2}; {4, 3}; {4, 4} ;

{5, 1}; {5, 2}; {5, 3}; {5, 4} ;

{6, 1}; {6, 2}; {6, 3}; {6, 4}

a) A = { sum of 5 is rolled}

therefore A consist of {1, 4}; { 2,3}, {3, 2}; {4,1}

Therefore P(A) = 4/24 = 1/6 = 0.166667

b) B = {sum of 5 or sum of 9 is rolled}.

So B = { {1, 4}; { 2,3}, {3, 2}; {4,1}; {5, 4}; {6, 3} }

Therefore P( B ) = 6/24 = 1/4 = 0.25

c) P( A and B ) = P( Common sample points in A and B ) = 4/24 = 1/6 = 0.166667